OpenStudy (anonymous):

6.78j-5.2=4.33j+2.15 j =3 how is this?

OpenStudy (anonymous):

collect like terms 6.78j-4.33j=2.15+5.2

OpenStudy (anonymous):

ahh, thanks.

OpenStudy (anonymous):

think about the pieces not the whole. put similar pieces together.

OpenStudy (anonymous):

yeah, we learned what like terms were.

OpenStudy (anonymous):

but what do you do after combining like terms?

OpenStudy (anonymous):

isolate the variable. 2.45j=2.45*j so if you divide both sides by 2.45, j is now alone

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

but how would you do 3.2k-4.3=12.6k+14.5?

OpenStudy (anonymous):

gather like terms

OpenStudy (anonymous):

so --- 3.2k+12.6k=4.3+14.5?